OpenStudy (anonymous):

I have to rationalize a radical again. -please wait for question to be posted-

OpenStudy (anonymous):

\[2\sqrt108\]

OpenStudy (turingtest):

put 108 into its prime factorization

OpenStudy (anonymous):

Which would be 3?

OpenStudy (turingtest):

not exactly

OpenStudy (anonymous):

:/

OpenStudy (turingtest):

prime factorization works like so: 108 since it's even we can divide it by 2 (or "factor out" a 2) so we get \[108=2\times54\]now since 54 is even we can factor out another 2 to get\[108=2\times2\times27=2^2\times27\]now continue in this way.... what number can we divide 27 by?

OpenStudy (anonymous):

Well, 3.

OpenStudy (anonymous):

OH WAIT 9

OpenStudy (turingtest):

well, you're right both times actually, since 9 is divisible by 3 so anything divisible by 9 is also divisible by 3 make sense?

OpenStudy (anonymous):

Yes. I was trying to answer the GCF. :P

OpenStudy (turingtest):

sorry I missed a 3 myself :/ let me do that again

OpenStudy (turingtest):

\[\sqrt{108}=\sqrt{2^2\times3\times9}=\sqrt{2^2\times3\times3^2}=(2\times3)\sqrt3\]since 27=3^3, and we can only pull out 3^2, we have to leave a 3 inside the radical

OpenStudy (turingtest):

so then what is\[2\sqrt{108}\]?

OpenStudy (anonymous):

So 6sqrt3? :D

OpenStudy (turingtest):

I just showed you the simplification of \[\sqrt{108}\]I think it should be your job to tell me how the 2 in front changes things

OpenStudy (turingtest):

hint:\[2\sqrt{108}=2\times(\sqrt{108})\]

OpenStudy (anonymous):

That's what I thought I just said. :|

OpenStudy (anonymous):

Was \[6\sqrt3\] wrong? :|

OpenStudy (anonymous):

12 root 3